two cars are moving with contant velocities of 15m/s and 25m/s if they are moving towards eachother and meet after 5 seconds find their initial interval
1
Expert's answer
2019-10-28T11:37:24-0400

The relative speed is


"v=15+25=40\frac{m}{s}"

Their initial interval:


"s=vt=(40)(5)=200\ m"